4.25x -12.5 ≤ 20.5
1
Expert's answer
2019-12-13T09:35:49-0500

"4.25x-12.5 \leq 20.5"

"4.25x \leq 12.5+20.5"

"4.25x \leq 33"

"x \leq \frac{33}{4.25}"

"x \leq 7.76471"